OpenStudy (anonymous):

Find (f ○ g)(x) for the indicated functions. f(x)=2x-6, g(x)= 3/x+6

OpenStudy (anonymous):

I believe you just plug the function f(x) into g(x) wereever you see x

OpenStudy (anonymous):

(f ○ g)(x) = 2 ( 3/x + 6) - 6 = 6/x + 6